OpenStudy (anonymous):

What are the zeroes of f(x) = x2 - 2x - 3? x = -3, 1 x = -3, -1 x = 3, 1 x = 3, -1

OpenStudy (anonymous):

Please help

OpenStudy (lυἶცἶ0210):

All you have to do is factor it and set it equal to zero. Do you know how to factor it?

OpenStudy (anonymous):

no i dont

OpenStudy (lυἶცἶ0210):

Well, what multiplies to -3 but adds up to -2?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

wait its A

OpenStudy (lυἶცἶ0210):

Nope, that would us \(x^2+2x-3\)

OpenStudy (anonymous):

oh okay umm..

OpenStudy (anonymous):

unless its B then

OpenStudy (lυἶცἶ0210):

If you can't factor, try using the quadratic formula \(\large \frac{-b \pm \sqrt{b^2-4ac}}{2a} \)

OpenStudy (anonymous):

ugh idk

OpenStudy (lυἶცἶ0210):

\(\Large \frac{-\color{blue}{b} \pm \sqrt{\color{blue}{b}^2-4\color{red}{a}* \color{green}{c}}}{2\color{red}{a}} \) \(\Large \color{red}{1}x^2\color{blue}{-2}x\color{green}{-3} \) Plug in the values

OpenStudy (anonymous):

D!?